# Imperative Learning: A Self-supervised Neural-Symbolic Learning Framework for Robot Autonomy

### v1

@@ 4.3 Constrained Optimization
$L({\bm{\mu}}_{k})=\sum_{k=0}^{T-1}\left(({\bm{\mu}}_{k}-{\bm{\mu}}_{k}^{\text{ref}})^{\top}\bm{Q}_{k}({\bm{\mu}}_{k}-{\bm{\mu}}_{k}^{\text{ref}})+\bm{p}_{k}({\bm{\mu}}_{k}-{\bm{\mu}}_{k}^{\text{ref}})\right),$ (56)
where ${\bm{\mu}}_{k}^{\text{ref}}$ is the reference state trajectory, and $\bm{Q}_{k}$ and $\bm{q}_{k}$ are the weight matrices to balance tracking performance and control effort, respectively. The optimal control action $\bm{u}_{k}$ will be commanded to both the modeled UAV (dynamics) and the real UAV, leading to the next state $\bm{x}_{k+1}$ and $\bm{x}^{I}_{k+1}$ (measured by the IMU network). The UL cost is defined as the Euclidean distance between $\bm{x}_{k+1}$ and $\bm{x}^{I}_{k+1}$
$U({\bm{\theta}})=\|\bm{x}_{k+1}^{I}-\bm{x}_{k+1}\|_{2}.$ (57)
@@ 4.5 Discrete Optimization
To overcome the above limitation, we reformulate MTSP as a BLO following the IL framework based on the control variate gradient estimator, as shown in Figure 8. This comprises a UL optimization, aiming to train an allocation network and a gradient estimator (surrogate network), and an LL optimization that solves decomposed single-agent TSPs. Specifically, the allocation network $f(\bm{x};{\bm{\theta}})$ takes the MTSP graph data $\bm{x}$ and produces a probability matrix, where its $(i,j)$-th entry represents the probability that city $i$ is assigned to agent $j$. This decomposes the MTSP into multiple single-agent TSPs by random sampling the city allocation $a_{j}\sim f(\bm{x};{\bm{\theta}})$ for agent $j$. Then, a non-differentiable TSP solver $g$ is employed to quickly find an optimal tour ${\bm{\mu}}$ for single-agent TSPs based on the allocation $a_{j}$, where $j=1,\cdots,M$. During training, the surrogate network $s({\bm{\gamma}})$, which serves as the memory module in the IL framework, is trained to have a low-variance gradient estimation through the non-differentiable reasoning engine $g$. This results in a self-supervised framework for solving MTSP, which we refer to as imperative MTSP (iMTSP) and formulate as:
@@
where $D$ computes the traveling distance given a tour, and $L$ returns the maximum route length among all agents. The UL optimization consists of two separate components: the UL cost $U_{\bm{\theta}}$ is to optimize the allocation network $f({\bm{\theta}})$, while $U_{\bm{\gamma}}$ is to optimize the surrogate network $s({\bm{\gamma}})$. Inspired by the control variant gradient estimator (67) and to have a lower gradient variance, we design the UL costs $U_{\bm{\theta}}$ and $U_{\bm{\gamma}}$ as
@@ 5.1.2 Datasets
To evaluate iA in diverse environments, we select three widely used datasets in path planning, Motion Planning (MP) (Bhardwaj et al., 2017), Matterport3D (Chang et al., 2017), and Maze (Ivanitskiy et al., 2023). Specifically, the MP dataset contains a set of binary obstacle maps from eight environments; the Matterport3D dataset provides diverse indoor truncated signed distance function (TSDF) maps; and the Maze dataset contains sophisticated mazes-type maps. To better evaluate the generalization ability, we randomly sample the start and goal positions following (Yonetani et al., 2021b) to provide more diverse maps.
@@ 5.1.5 Qualitative Performance
Figure 10: Real-world experiment for local path planning using iPlanner with a legged robot. The red curve indicates the robot’s trajectory starting from right to left, which begins inside a building and then navigates to the outdoors. The robot follows a series of waypoints (blue) and plans in different scenarios marked by green boxes including (A) passing through doorways, (B, D, E) circumventing both static and dynamic obstacles, and (B, F) ascending and descending stairs.
@@ 5.3.4 Metrics
5.3.5 Robustness to Initial Conditions
@@ 5.4.6 Effectiveness Validation
Figure 16: Comparison of the VO trajectories at different imperative iterations in iSLAM alongside the ground truth trajectories.
5.4.7 Adaptation to New Environments
@@ 5.5.1 Baselines
To thoroughly evaluate the performance of different methods for MTSP, we compare our iMTSP with both classic methods and learning-based methods. This includes a well-known meta-heuristic method implemented by the Google OR-Tools routing library (Perron and Didier, 2023) and an RL-based method (Hu et al., 2020), which will be referred to as Google OR-Tools and RL baseline for abbreviation, respectively. Specifically, Google OR-Tools can provide near-optimal solutions [^1] Google OR-Tools is a meta-heuristic and cannot guarantee the optimality of solutions. However, it is very efficient for small-scale problem and wildly used for academic and commercial purposes. Thus, in our experiments, it can be seen as an “ideal” solver providing “optimal” solutions. to MTSPs with a few hundred cities, while the RL baseline can solve larger-scale MTSPs such as with one thousand cities. We cannot compare with other methods such as (Liang et al., 2023; Gao et al., 2023; Park et al., 2021) because they either don’t provide source code or cannot be applied to our settings.
@@ 5.5.2 Datasets
Due to the uniqueness of MTSP, we cannot deploy MTSP algorithms on such a large-scale robot team, thus we build a simulation for comparison. Specifically, all the data instances are generated by uniformly sampling points in a unit rectangular so that both $x$ and $y$ coordinates are in the range of 00 to $1$. The test set consists of i.i.d. samples but with a larger number ($400$ to $1000$) of cities as aforementioned.
Table 16: Performance of different methods on large-scale MTSPs. iMTSP significantly outperforms the Google OR-Tools (Perron and Didier, 2023) and provides an average of $3.2\pm 0.01\%$ shorter maximum route length than the RL baseline (Hu et al., 2020).
@@ 5.5.3 Implementation Details
The structure of our allocation network is similar to that in (Hu et al., 2020), where the cities and agents are embedded into $64$ dimensional vectors. The surrogate network is a three-layer MLP whose hidden dimension is $256$ and whose activation function is “tanh”. Additionally, Google OR-Tools provides the TSP solver in our iMTSP. It uses the “Global Cheapest Arc” strategy to generate its initial solutions and uses “Guided local search” for local improvements. We use the best-recommended settings in Google OR Tools for all experiments.
@@ 5.5.4 Quantitative Performance
Table 17: Computing time of iMTSP dealing with $1000$ cities. Due to a similar network structure, the RL baseline (Hu et al., 2020) has roughly the same runtime efficiency compared with ours, while Google OR-Tools (Perron and Didier, 2023) cannot find feasible solutions within a given time budget.

### v2

@@ 4.3 Constrained Optimization
$L({\bm{\mu}}_{k})\doteq\sum_{k=0}^{T-1}\left(\Delta\bm{\mu}_{k}^{\top}\bm{Q}_{k}\Delta\bm{\mu}_{k}+\bm{p}_{k}\Delta\bm{\mu}_{k}\right),$ (56)
where $\Delta\bm{\mu}_{k}={\bm{\mu}}_{k}-{\bm{\mu}}_{k}^{\text{ref}}$, ${\bm{\mu}}_{k}^{\text{ref}}$ is the reference state trajectory, and $\bm{Q}_{k}$ and $\bm{q}_{k}$ are the weight matrices to balance tracking performance and control effort, respectively. The optimal control action $\bm{u}_{k}$ will be commanded to both the modeled UAV (dynamics) and the real UAV, leading to the next state $\bm{x}_{k+1}$ and $\bm{x}^{I}_{k+1}$ (measured by the IMU network). The UL cost is defined as the Euclidean distance between $\bm{x}_{k+1}$ and $\bm{x}^{I}_{k+1}$
$U({\bm{\theta}})\doteq\|\bm{x}_{k+1}^{I}-\bm{x}_{k+1}\|_{2}.$ (57)
@@ 4.5 Discrete Optimization
To overcome the above limitation, we reformulate MTSP as a BLO following the IL framework based on the control variate gradient estimator, as shown in Figure 8. This comprises a UL optimization, aiming to train an allocation network and a surrogate network assisting gradient estimation, and an LL optimization that solves decomposed single-agent TSPs. Specifically, the allocation network $f(\bm{x};{\bm{\theta}})$ takes the MTSP graph data $\bm{x}$ and produces a probability matrix, where its $(i,j)$-th entry represents the probability that city $i$ is assigned to agent $j$. This decomposes the MTSP into multiple single-agent TSPs by random sampling the city allocation $a_{j}\sim f(\bm{x};{\bm{\theta}})$ for agent $j$. Then, a non-differentiable TSP solver $g$ is employed to quickly find an optimal tour ${\bm{\mu}}$ for single-agent TSPs based on the allocation $a_{j}$, where $j=1,\cdots,M$. During training, the surrogate network $s({\bm{\gamma}})$, which serves as the memory module in the IL framework, is trained to have a low-variance gradient estimation through the non-differentiable reasoning engine $g$. This results in a self-supervised formulation for MTSP, which we refer to as imperative MTSP (iMTSP):
@@
where $D$ computes the traveling cost given a tour, and $L$ returns the maximum route length among all agents. The UL optimization consists of two separate components: the UL cost $U_{\bm{\theta}}$ is to optimize the allocation network $f({\bm{\theta}})$, while $U_{\bm{\gamma}}$ is to optimize the surrogate network $s({\bm{\gamma}})$. Inspired by the control variant gradient estimator (67) and to have a lower gradient variance, we design the UL costs $U_{\bm{\theta}}$ and $U_{\bm{\gamma}}$ as
@@ 5.1.2 Datasets
To evaluate iA in diverse environments, we select three widely used datasets in path planning, Motion Planning (MP) (Bhardwaj et al., 2017), Matterport3D (Chang et al., 2017), and Maze (Ivanitskiy et al., 2023). Specifically, the MP dataset contains a set of binary obstacle maps from eight environments; the Matterport3D dataset provides diverse indoor truncated signed distance function (TSDF) maps; and the Maze dataset contains sophisticated mazes-type maps. To better evaluate the generalization ability, we randomly sample the start and goal positions following (Yonetani et al., 2021b) to provide more diverse maps. It is essential to note that all methods only use the MP dataset for training, and use Matterport3D and Maze for testing, which is to demonstrate their generalization abilities.
@@ 5.1.5 Qualitative Performance
Figure 10: Real-world experiment for local path planning using iPlanner with a legged robot. The red curve indicates the robot’s trajectory from right to left, beginning inside a building and then navigating to the outdoors. The robot follows a series of waypoints (blue) and plans in different scenarios marked by green boxes including (A) passing through doorways, (B, D, E) circumventing both static and dynamic obstacles, and (B, F) ascending and descending stairs.
@@ 5.3.1 Baselines
5.3.2 Implementation Details
5.3.3 Simulation Environment
5.3.4 Metrics
@@ 5.4.3 Accuracy Evaluation
5.4.4 Robustness Assessment
Figure 16: The predicted trajectories from the front-end are improved concerning the number of imperative iterations in iSLAM.
@@ 5.5.1 Baselines
To thoroughly evaluate the performance of different methods for MTSP, we compare our iMTSP with both classic methods and learning-based methods. This includes a well-known meta-heuristic method implemented by the Google OR-Tools routing library (Furnon and Perron, 2023) and an RL-based method (Hu et al., 2020), which will be referred to as Google OR-Tools and RL baseline for abbreviation, respectively. Specifically, Google OR-Tools can provide near-optimal solutions [^3] Google OR-Tools is a meta-heuristic and cannot guarantee the optimality of solutions. However, it is efficient for small-scale problems and wildly used for academic and commercial purposes. Thus, in our experiments, it can be seen as an “ideal” solver providing “optimal” solutions. to MTSPs with a few hundred cities, while the RL baseline can solve larger-scale MTSPs such as with one thousand cities. We cannot compare with other methods such as (Liang et al., 2023; Gao et al., 2023; Park et al., 2021) because they either don’t provide source code or cannot be applied to our settings.
@@ 5.5.2 Datasets
Due to the uniqueness of MTSP, we cannot deploy MTSP algorithms on such a large-scale robot team, thus we build a simulation for comparison. Specifically, all the data instances are generated by uniformly sampling points in a unit rectangular so that both $x$ and $y$ coordinates are in the range of 00 to $1$. The test set consists of i.i.d. samples but with a larger number ($400$ to $1000$) of cities as aforementioned. Note that the proposed approach is applicable to general graphs with arbitrary costs assigned to edges, although our dataset is constrained in the 2-D space.
Table 16: Performance of different methods on large-scale MTSPs. iMTSP significantly outperforms the Google OR-Tools (Furnon and Perron, 2023) and provides an average of $3.2\pm 0.01\%$ shorter maximum route length than the RL baseline (Hu et al., 2020).
@@ 5.5.3 Implementation Details
The structure of our allocation network is similar to that in (Hu et al., 2020), where the cities and agents are embedded into $64$ dimensional vectors. The surrogate network is a three-layer MLP whose hidden dimension is $256$ and whose activation function is “tanh”. Additionally, Google OR-Tools provides the TSP solver in our iMTSP. It uses the “Global Cheapest Arc” strategy to generate its initial solutions and uses “Guided local search” for local improvements. We use the best-recommended settings in Google OR Tools for all experiments. All tests were conducted locally using the same desktop-level computer.
@@ 5.5.4 Quantitative Performance
Table 17: Computing time of iMTSP dealing with $1000$ cities. Due to a similar network structure, the RL baseline (Hu et al., 2020) has roughly the same runtime efficiency compared with ours, while Google OR-Tools (Furnon and Perron, 2023) cannot find feasible solutions within a given time budget.